# Improving Bilinear RNNs with Closed-loop Control

### v4

@@ Abstract
footnotetext: RCorresponding authors (yuxliang@outlook.com, sunweigao@outlook.com). * Work done during Yongqi Pan’s internship at Hong Kong University of Science and Technology (Guangzhou).
@@ 1 Introduction
Autoregressive Transformers [95] have become a foundation of modern AI, primarily due to the efficient parallel computation made possible by softmax-based self-attention. This mechanism enables effective memory scaling by directly appending key and value vectors into the KV cache, which contributes to strong performance on tasks such as in-context learning and long-context retrieval. However, this design also comes with challenges, including quadratic time complexity and unbounded memory growth during inference [59], which limits the model’s scalability for long-sequence tasks. To address these, numerous improvements have been introduced, including sliding window attention [8, 25], sparse attention techniques [64, 111, 103], and efficient KV cache management [59].
@@
Meanwhile, efficient sequence-mixing approaches such as gated linear attention [112, 79, 22, 107, 105, 5] and selective state space models (SSMs) [38, 26, 84] offer a compelling alternative. These models aim to establish a linear key-value associative memory [105, 34] register with constant states and data-(in)dependent gating ($\bm{\alpha}$, $\bm{\beta}$), as presented by $\bm{S}_{t}=\bm{\alpha}_{(t)}\bm{S}_{t-1}+\bm{\beta}_{(t)}\bm{v}_{t}\bm{k}_{t}^{\intercal}$. The inherent recurrent structure of these models enables them to maintain constant memory overhead and $\mathcal{O}(1)$ time complexity during inference. Despite these models originating from distinct theoretical frameworks, for example, early linear attentions [53, 82] like Linformer [98] attempt to reformulate quadratic attention computation as $\bm{O}=\phi(\bm{Q})(\phi(\bm{K})^{\intercal}\bm{V})$ with kernel mapping $\phi$, while original SSMs [39, 42] like S4 [41] intent to parameterize a continuous dynamical system to a discrete form, recent literature [22, 7, 52, 6] have unified these models to the concept of Linear RNNs.
@@
The data-dependent gating in Linear RNNs provides a dynamic memory management similar to the adaptive, weighted information fusion in softmax attention. This allows such models to selectively update and retain relevant information, enabling Linear RNNs to serve as practical replacements for Transformers in many downstream tasks [48, 62, 49]. However, this mechanism remains heuristic; that is, the model lacks a criterion for determining which memories to forget, and all key-value associations are forgotten uniformly, rendering the process less targeted and efficient [105].
@@
To address this, recent models such as DeltaNet [108, 105], RWKV-7 [73], and TTT [90] have advanced state transition to generalized Householder transformations [51], enhancing model’s learning capacity and enabling supervised memory control via the Delta learning rule [101]. These architectural shifts foster richer interactions between the internal state $\bm{S}$ and the input information $\bm{k}$, moving beyond simple linear key-value memory registers and resembling bilinear dynamics. Consequently, we refer to these models as Bilinear RNNs in this paper. Further improvements [7, 52, 6] have built upon TTT by introducing higher-order nonlinear optimization or MLP-based deep memory, giving rise to modern Nonlinear RNNs, which enhance expressiveness but sacrifice the ability to perform chunk-wise parallelism over the sequence. In summary, research in this field is still in its early stages, and there remain open challenges in achieving a good balance between model expressiveness and hardware-efficient implementation during pretraining.
@@
Table 1: Memorizing Mechanismes in state-of-the-art Sequence Modeling Methods. Softmax Attention ensures precise memory storage, while sliding window attention constrains storage space. Linear RNNs (lines 3-6) use data-(in)dependent gate for unsupervised memory management, with Mamba2 approximating the forget gate $\alpha$ to 1, forcing the model to forget. Bilinear RNNs (lines 7-9) is no longer a linear key-value memory register $\bm{S}_{t+1}=(\bm{\alpha},\bm{\beta})@{(\bm{S}_{t},\bm{k}_{t}^{\intercal}\bm{v}_{t})}^{\intercal}$, which supervise the management process based on the Delta learning rule – effectively equivalent to minimizing a Stochastic Gradient Descent $\nabla_{S}\left\|\bm{S}_{t}\bm{k}_{t}-\bm{v}_{t}\right\|^{2}$. When ignoring layer normalization and residual components, TTT-Linear can also be categorized as a bilinear RNN. As modern Nonlinear RNNs, MIRAS and its variants (e.g., TTT-MLP, Titans [7], Lattice [52]) have stronger expressiveness because of the nonlinearity $g$, high-order optimizations, and MLP-based deep memory, but are limited by the chunk-wise parallelism (§3). Our proposed Comba further improves Bilinear RNNs by closed-loop control.
@@
Model Memorizing with Gate Optimization Objective $\mathcal{L}_{\text{min}}$ Softmax Attention: Conductivist-based infinite key-value associative memory registers SA [95] $\bm{S}_{t}=\bm{S}_{t-1}.\operatorname{append}(\bm{k}_{t},\bm{v}_{t})$ $\sum_{i=1}^{t}\exp(\bm{q}_{t}^{\intercal}\bm{k}_{i})\left\|\bm{v}-\bm{v}_{i}\right\|^{2}$ [97] SWA [8] $\bm{S}_{t}=\bm{S}_{t-1}.\operatorname{append}(\bm{k}_{t},\bm{v}_{t}).\operatorname{drop}(\bm{k}_{t-M},\bm{v}_{t-M})$ $\sum_{i=t-M}^{t}\exp(\bm{q}_{t}^{\intercal}\bm{k}_{i})\left\|\bm{v}-\bm{v}_{i}\right\|^{2}$ Linear RNNs: Inductivist-based finite key-value associative memory registers LA [98] $\bm{S}_{t}=\bm{S}_{t-1}+\bm{v}_{t}\bm{k}_{t}^{\intercal}$ $\left\langle\bm{S}_{t}\bm{k}_{t},\bm{v}_{t}\right\rangle$ GLA [107] $\bm{S}_{t}=\bm{S}_{t-1}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\operatorname{diag}(\bm{\alpha}_{t})}+\bm{v}_{t}\bm{k}_{t}^{\intercal}$ $\left\langle\bm{S}_{t}\bm{k}_{t},\bm{v}_{t}\right\rangle+\frac{\operatorname{Diag}(\bm{1}-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bm{\alpha}_{t}})}{2}\left\|\bm{S}_{t}\right\|_{2}^{2}$ HGRN2 [79] $\bm{S}_{t}=\bm{S}_{t-1}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\operatorname{diag}(\bm{\alpha}_{t})}+\bm{v}_{t}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(\bm{1-\alpha}_{t}})^{\intercal}$ $\left\langle\bm{S}_{t}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(\bm{1-\alpha}_{t}}),\bm{v}_{t}\right\rangle+\frac{\operatorname{Diag}(\bm{1}-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bm{\alpha}_{t}})}{2}\left\|\bm{S}_{t}\right\|_{2}^{2}$ Mamba2 [26] $\bm{S}_{t}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\alpha_{t}^{\sim 1}}\bm{S}_{t-1}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\beta_{t}^{\sim 0}}\bm{v}_{t}\bm{k}_{t}^{\intercal}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\beta_{t}}\left\langle\bm{S}_{t}\bm{k}_{t},\bm{v}_{t}\right\rangle+\frac{1-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\alpha_{t}}}{2}\left\|\bm{S}_{t}\right\|_{2}^{2}$ Bilinear RNNs: Moving beyond linear key-value memory registers with memory correction G-DeltaNet [105] $\bm{S}_{t}=\bm{S}_{t-1}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\alpha_{t}^{\sim 1}\left(\bm{I}_{t}-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\beta}_{t}\bm{k}_{t}\bm{k}_{t}^{\intercal}\right)}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\beta}_{t}\bm{v}_{t}\bm{k}_{t}^{\intercal}$ $\frac{1}{2}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\beta_{t}}\left\|\frac{1}{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\alpha_{t}}}\bm{v}_{t}-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\alpha_{t}}\bm{S}_{t}\bm{k}_{t}\right\|^{2}+\frac{\bm{1}-\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\alpha_{t}}{2}\left\|\bm{S}_{t}\right\|_{2}^{2}$ RWKV7 [73] $\bm{S}_{t}=\bm{S}_{t-1}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(\operatorname{diag}(\bm{\alpha}_{t})-\beta_{t}\hat{\bm{k}}_{t}\hat{\bm{k}}_{t}^{\intercal})}+\bm{v}_{t}\tilde{\bm{k}}_{t}^{\intercal}$ $\frac{1}{2}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\beta_{t}}\left\|\frac{1}{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\beta_{t}}}\bm{v}_{t}-\bm{S}_{t}\bm{k}_{t}\right\|^{2}+\frac{\operatorname{Diag}(\bm{1}-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bm{\alpha}_{t}})}{2}\left\|\bm{S}_{t}\right\|_{2}^{2}$ Comba (ours) $\bm{S}_{t}=\bm{S}_{t-1}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\left(\alpha_{t}^{\sim 1}-\beta_{t}^{\downarrow}\bm{k}_{t}\bm{k}_{t}^{\intercal}\right)}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\beta_{t}^{{}^{\uparrow}}}\bm{v}_{t}\bm{k}_{t}^{\intercal}$ $\frac{1}{2}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\beta_{t}}\left\|\bm{v}_{t}-\bm{S}_{t}\bm{k}_{t}\right\|^{2}+\frac{\bm{1}-\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\alpha_{t}}{2}\left\|\bm{S}_{t}\right\|_{2}^{2}-\left\langle\bm{q}_{t},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}d}\bm{k}_{t}\right\rangle$ (Modern) Nonlinear RNNs: Stronger expressiveness and memory capacity, but limited in chunk-wise parallelism. TTT-MLP [90] $\bm{S}_{t}(\cdot)=\bm{S}_{t-B}(\cdot)-\sum_{i=1}^{B}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\beta_{i}}\nabla_{S}\left\|\psi(\bm{S}_{t-B}(\bm{k}_{i}))-\bm{v}_{i}\right\|^{2}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\beta_{i}}\left\|\bm{v}_{i}-\psi(\bm{S}_{j}(\bm{k}_{i}))\right\|^{2}$ MIRAS [6] $\bm{S}_{t}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\alpha_{t}}\bm{S}_{t-1}-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\beta_{t}}\nabla_{S}\left\|g(\psi(\bm{S}_{t-1}),\bm{k}_{t})-\bm{v}_{t}\right\|_{p}^{p}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\beta_{t}}\left\|\bm{v}_{t}-g(\psi(\bm{S}_{t}),\bm{k}_{t})\right\|_{p}^{p}+\frac{\operatorname{Diag}(\bm{1}-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bm{\alpha}_{t}})}{2}\left\|\bm{S}_{t}\right\|_{2}^{2}$ $\bm{S}_{t}$ denotes memory, while $\bm{k}_{t},\bm{v}_{t}$ are key-value pairs. $\alpha_{t}^{\sim 1}$ denotes close to 1. $\beta_{t}^{\downarrow},\beta_{t}^{\uparrow}$ represent smaller/bigger factors. Early Nonlinear RNNs, such as LSTM [46] and GRU [21] are omitted on the table. Softmax attention can also be interpreted as an L1 loss with a kernel function [114].
@@
  • Inspired by closed-loop control theory, we propose a novel Bilinear RNN architecture named Comba. Unlike previous models, Comba features a scalar-plus-low-rank (SPLR) state transition and applies feedback control to the query vector during the output. We further develop a hardware-friendly implementation of Comba using chunk-wise parallelism in Triton [93], which achieves a 40% speed improvement in forward propagation compared to Gated-DeltaNet. (Section 3)

### v5

@@ Abstract
footnotetext: RCorresponding authors (yuxliang@outlook.com, sunweigao@outlook.com). J Hu (jhu110@connect.hkust-gz.edu.cn). * Work done during Yongqi Pan’s internship at HKUST(GZ).
@@ 1 Introduction
Autoregressive Transformers [vaswani2017attention] have become a foundation of modern AI, primarily due to the efficient parallel computation made possible by softmax-based self-attention. This mechanism enables effective memory scaling by directly appending key and value vectors into the KV cache, which contributes to strong performance on tasks such as in-context learning and long-context retrieval. However, this design also comes with challenges, including quadratic time complexity and unbounded memory growth during inference [li2024survey], which limits the model’s scalability for long-sequence tasks. To address these, numerous improvements have been introduced, including sliding window attention [beltagy2020longformer, dao2022flashattention], sparse attention techniques [lu2025moba, yuan2025native, xu2025xattention], and efficient KV cache management [li2024survey].
@@
Meanwhile, efficient sequence-mixing approaches such as gated linear attention [zhang2024gated, qin2024hgrn2, chou2024metala, yang2023gated, yang2024gated, beck2024xlstm] and selective state space models (SSMs) [gu2023mamba, dao2024transformers, smith2022simplified] offer a compelling alternative. These models aim to establish a linear key-value associative memory [yang2024gated, gershman2025key] register with constant states and data-(in)dependent gating ($\bm{\alpha}$, $\bm{\beta}$), as presented by $\bm{S}_{t}=\bm{\alpha}_{(t)}\bm{S}_{t-1}+\bm{\beta}_{(t)}\bm{v}_{t}\bm{k}_{t}^{\intercal}$. The inherent recurrent structure of these models enables them to maintain constant memory overhead and $\mathcal{O}(1)$ time complexity during inference. Despite these models originating from distinct theoretical frameworks, for example, early linear attentions [katharopoulos2020transformers, shen2021efficient] like Linformer [wang2020linformer] attempt to reformulate quadratic attention computation as $\bm{O}=\phi(\bm{Q})(\phi(\bm{K})^{\intercal}\bm{V})$ with kernel mapping $\phi$, while original SSMs [gu2020hippo, gu2022train] like S4 [gu2021efficiently] intent to parameterize a continuous dynamical system to a discrete form, recent literature [chou2024metala, behrouz2024titans, karami2025lattice, behrouz2025s] have unified these models to the concept of Linear RNNs.
@@
The data-dependent gating in Linear RNNs provides a dynamic memory management similar to the adaptive, weighted information fusion in softmax attention. This allows such models to selectively update and retain relevant information, enabling Linear RNNs to serve as practical replacements for Transformers in many downstream tasks [hu2024time, liu2024vmamba, hu2024state]. However, this mechanism remains heuristic; that is, the model lacks a criterion for determining which memories to forget, and all key-value associations are forgotten uniformly, rendering the process less targeted and efficient [yang2024gated].
@@
To address this, recent models such as DeltaNet [yang2024parallelizing, yang2024gated], RWKV-7 [peng2025rwkv], and TTT [sun2024learning] have advanced state transition to generalized Householder transformations [joffrain2006accumulating], enhancing model’s learning capacity and enabling supervised memory control via the Delta learning rule [widrow1988adaptive]. These architectural shifts foster richer interactions between the internal state $\bm{S}$ and the input information $\bm{k}$, moving beyond simple linear key-value memory registers and resembling bilinear dynamics. Consequently, we refer to these models as Bilinear RNNs in this paper. Further improvements [behrouz2024titans, karami2025lattice, behrouz2025s] have built upon TTT by introducing higher-order nonlinear optimization or MLP-based deep memory, giving rise to modern Nonlinear RNNs, which enhance expressiveness but sacrifice the ability to perform chunk-wise parallelism over the sequence. In summary, research in this field is still in its early stages, and there remain open challenges in achieving a good balance between model expressiveness and hardware-efficient implementation during pretraining.
@@
Table 1: Memorizing Mechanismes in state-of-the-art Sequence Modeling Methods. Softmax Attention ensures precise memory storage, while sliding window attention constrains storage space. Linear RNNs (lines 3-6) use data-(in)dependent gate for unsupervised memory management, with Mamba2 approximating the forget gate $\alpha$ to 1, forcing the model to forget. Bilinear RNNs (lines 7-9) is no longer a linear key-value memory register $\bm{S}_{t+1}=(\bm{\alpha},\bm{\beta})@{(\bm{S}_{t},\bm{k}_{t}^{\intercal}\bm{v}_{t})}^{\intercal}$, which supervise the management process based on the Delta learning rule – effectively equivalent to minimizing a Stochastic Gradient Descent $\nabla_{S}\left\|\bm{S}_{t}\bm{k}_{t}-\bm{v}_{t}\right\|^{2}$. When ignoring layer normalization and residual components, TTT-Linear can also be categorized as a bilinear RNN. As modern Nonlinear RNNs, MIRAS and its variants (e.g., TTT-MLP, Titans [behrouz2024titans], Lattice [karami2025lattice]) have stronger expressiveness because of the nonlinearity $g$, high-order optimizations, and MLP-based deep memory, but are limited by the chunk-wise parallelism (§3). Our proposed Comba further improves Bilinear RNNs by closed-loop control.
@@
Model Memorizing with Gate Optimization Objective $\mathcal{L}$ Softmax Attention: Conductivist-based infinite key-value associative memory registers SA [vaswani2017attention] $\bm{S}_{t}=\bm{S}_{t-1}.\operatorname{append}(\bm{k}_{t},\bm{v}_{t})$ $\sum_{i=1}^{t}\exp(\bm{q}_{t}^{\intercal}\bm{k}_{i})\left\|\bm{v}-\bm{v}_{i}\right\|^{2}$ [wang2025test] SWA [beltagy2020longformer] $\bm{S}_{t}=\bm{S}_{t-1}.\operatorname{append}(\bm{k}_{t},\bm{v}_{t}).\operatorname{drop}(\bm{k}_{t-M},\bm{v}_{t-M})$ $\sum_{i=t-M}^{t}\exp(\bm{q}_{t}^{\intercal}\bm{k}_{i})\left\|\bm{v}-\bm{v}_{i}\right\|^{2}$ Linear RNNs: Inductivist-based finite key-value associative memory registers LA [wang2020linformer] $\bm{S}_{t}=\bm{S}_{t-1}+\bm{v}_{t}\bm{k}_{t}^{\intercal}$ $\left\langle\bm{S}_{t}\bm{k}_{t},\bm{v}_{t}\right\rangle$ GLA [yang2023gated] $\bm{S}_{t}=\bm{S}_{t-1}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\operatorname{diag}(\bm{\alpha}_{t})}+\bm{v}_{t}\bm{k}_{t}^{\intercal}$ $\left\langle\bm{S}_{t}\bm{k}_{t},\bm{v}_{t}\right\rangle+\frac{1}{2}\left\|\sqrt{\operatorname{diag}(\bm{1}-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bm{\alpha}_{t}})}\bm{S}_{t}\right\|_{F}^{2}$ HGRN2 [qin2024hgrn2] $\bm{S}_{t}=\bm{S}_{t-1}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\operatorname{diag}(\bm{\alpha}_{t})}+\bm{v}_{t}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(\bm{1-\alpha}_{t}})^{\intercal}$ $\left\langle\bm{S}_{t}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(\bm{1-\alpha}_{t}}),\bm{v}_{t}\right\rangle+\frac{1}{2}\left\|\sqrt{\operatorname{diag}(\bm{1}-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bm{\alpha}_{t}})}\bm{S}_{t}\right\|_{2}^{2}$ Mamba2 [dao2024transformers] $\bm{S}_{t}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\alpha_{t}^{\sim 1}}\bm{S}_{t-1}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\beta_{t}^{\sim 0}}\bm{v}_{t}\bm{k}_{t}^{\intercal}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\beta_{t}}\left\langle\bm{S}_{t}\bm{k}_{t},\bm{v}_{t}\right\rangle+\frac{1}{2}\left\|\sqrt{1-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\alpha_{t}}}\bm{S}_{t}\right\|_{2}^{2}$ Bilinear RNNs: Moving beyond linear key-value memory registers with memory correction G-DeltaNet [yang2024gated] $\bm{S}_{t}=\bm{S}_{t-1}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\alpha_{t}^{\sim 1}\left(\bm{I}_{t}-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\beta}_{t}\bm{k}_{t}\bm{k}_{t}^{\intercal}\right)}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\beta}_{t}\bm{v}_{t}\bm{k}_{t}^{\intercal}$ $\frac{1}{2}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\alpha_{t}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\beta_{t}}\left\|\frac{1}{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\alpha_{t}}}\bm{v}_{t}-\bm{S}_{t}\bm{k}_{t}\right\|^{2}+\frac{1}{2}\left\|\sqrt{\bm{1}-\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\alpha_{t}}\bm{S}_{t}\right\|_{2}^{2}$ RWKV7 [peng2025rwkv] $\bm{S}_{t}=\bm{S}_{t-1}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(\operatorname{diag}(\bm{\alpha}_{t})-\beta_{t}\hat{\bm{k}}_{t}\hat{\bm{k}}_{t}^{\intercal})}+\bm{v}_{t}\tilde{\bm{k}}_{t}^{\intercal}$ $\frac{1}{2}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\beta_{t}}\left\|\frac{1}{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\beta_{t}}}\bm{v}_{t}-\bm{S}_{t}\bm{k}_{t}\right\|^{2}+\frac{1}{2}\left\|\sqrt{\operatorname{diag}(\bm{1}-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bm{\alpha}_{t}})}\bm{S}_{t}\right\|_{2}^{2}$ Comba (ours) $\bm{S}_{t}=\bm{S}_{t-1}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\left(\alpha_{t}^{\sim 1}-\beta_{t}^{\downarrow}\bm{k}_{t}\bm{k}_{t}^{\intercal}\right)}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\beta_{t}^{{}^{\uparrow}}}\bm{v}_{t}\bm{k}_{t}^{\intercal}$ $\frac{1}{2}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\beta_{t}}\left\|\bm{v}_{t}-\bm{S}_{t}\bm{k}_{t}\right\|^{2}+\frac{1}{2}\left\|\sqrt{\bm{1}-\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\alpha_{t}}\bm{S}_{t}\right\|_{2}^{2}-\left\langle\bm{q}_{t},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}d}\bm{k}_{t}\right\rangle$ (Modern) Nonlinear RNNs: Stronger expressiveness and memory capacity, but limited in chunk-wise parallelism. TTT-MLP [sun2024learning] $\bm{S}_{t}(\cdot)=\bm{S}_{t-B}(\cdot)-\sum_{i=1}^{B}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\beta_{i}}\nabla_{S}\left\|\psi(\bm{S}_{t-B}(\bm{k}_{i}))-\bm{v}_{i}\right\|^{2}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\beta_{i}}\left\|\bm{v}_{i}-\psi(\bm{S}_{j}(\bm{k}_{i}))\right\|^{2}$ MIRAS [behrouz2025s] $\bm{S}_{t}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\alpha_{t}}\bm{S}_{t-1}-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\beta_{t}}\nabla_{S}\left\|g(\psi(\bm{S}_{t-1}),\bm{k}_{t})-\bm{v}_{t}\right\|_{p}^{p}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\beta_{t}}\left\|\bm{v}_{t}-g(\psi(\bm{S}_{t}),\bm{k}_{t})\right\|_{p}^{p}+\frac{1}{2}\left\|\sqrt{\operatorname{diag}(\bm{1}-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bm{\alpha}_{t}})}\bm{S}_{t}\right\|_{2}^{2}$ $\bm{S}_{t}$ denotes memory, while $\bm{k}_{t},\bm{v}_{t}$ are key-value pairs. $\alpha_{t}^{\sim 1}$ denotes close to 1. $\beta_{t}^{\downarrow},\beta_{t}^{\uparrow}$ represent smaller/bigger factors. Early Nonlinear RNNs, such as LSTM [hochreiter1997long] and GRU [cho2014learning] are omitted on the table. Softmax attention can also be interpreted as an L1 loss with a kernel function [zhong2025understanding].
@@
  • Inspired by closed-loop control theory, we propose a novel Bilinear RNN architecture named Comba. Unlike previous models, Comba features a scalar-plus-low-rank (SPLR) state transition and applies feedback control to the query vector during the output. We further develop a hardware-friendly implementation of Comba using chunk-wise parallelism in Triton [tillet2019triton], which achieves a 40% speed improvement in forward propagation compared to Gated-DeltaNet. (Section 3)